OpenStudy (anonymous):

How would one differentiate the following: 2x+3/√(1-2x)

OpenStudy (anonymous):

You can use the quotient rule.

OpenStudy (anonymous):

Are you familiar with that?

OpenStudy (anonymous):

yes I am familiar with this

OpenStudy (anonymous):

so u = 2x+3 u' = 2 v = √(1-2x) v' = -1/√(1-2x)

OpenStudy (anonymous):

The quotient rule says dy/dx = (u'v - uv')/v^2

OpenStudy (anonymous):

ok and from there I begin to distribute, correct?

OpenStudy (anonymous):

u can do it with quotient rule

OpenStudy (anonymous):

Thats it follow (u'v - uv')/v^2

OpenStudy (anonymous):

 use this this is the division rule